# Formal Degrees and Parabolic Induction: the Maximal Generic Case

### v2

@@ 7. An Application to $\mathrm{G}_{2}$
In this section, as an application of the main theorem, we verify the formal degree conjecture for discrete series of split $\mathrm{G}_{2}$ supported on a maximal Levi subgroup.
Let $\mathbf{G}=\mathrm{G}_{2}$ be the split exceptional group of type $\mathrm{G}_{2}$ over $F$. We first recall the classification of discrete series representations of $G=\mathbf{G}(F)$ supported on a maximal Levi, following Muić [Mui97] (also cf. [GS23, §2]).
@@
For convenience of the reader, we list the related normalizations and the decomposition $r=\oplus_{i}r_{i}$ (the adjoint representation of ${{}^{L}\!M}$ on ${{}^{L}\mathfrak{n}}$) for these two cases as follows.
@@
    In this case the representation $r$ of ${}^{L}\!M_{\alpha}$ on ${}^{L}\!\mathfrak{n}_{\alpha}$ decomposes as
@@
    In this case the representation $r$ of ${}^{L}\!M_{\beta}$ on ${}^{L}\!\mathfrak{n}_{\beta}$ decomposes as
@@
The full explicit local Langlands correspondence for $\mathrm{G}_{2}$ has been established in [GS23a] based on a series work in exceptional theta correspondence. The conjectural L-parameters for non-supercuspidal discretes series have already been classified in [GS23, §3.5], which we now recall.
@@
$\pi$ $\varphi(\mathrm{SL}_{2})$ $\varphi(\mathrm{W}_{F})$ $\mathcal{S}_{\varphi}$
@@
by [GS23, §3.5]. (Note that the representations $\pi_{\alpha}(\tau,1/2)$, $\pi_{\beta}(\tau,1/2)$, $\pi_{\beta}(\tau,1)$ above are denoted by $\delta_{P}(\tau)$, $\delta_{Q}(\tau)$, $\pi_{\mathrm{gen}}[\tau]$, respectively, in [GS23].)
@@
Clearly, the orders of the component groups are compatible with the formal degree conjecture by (6.4). The compatibility of Langlands-Shahidi $\gamma$-factors and those from LLC for split $\mathrm{G}_{2}$ has been verified in [Sha89]. Of course, the formal degree conjecture is known for $\mathbf{M}_{\alpha},\mathbf{M}_{\beta}\simeq\mathrm{GL}_{2}$; thus as an application we have verified the formal degree conjecture for discrete series of $\mathrm{G}_{2}$ supported on maximal Levi subgroups.
@@ Acknowledgements
The author would like to first thank his advisor A. Ichino for suggesting this problem and many helpful discussions. He thanks professor W.T. Gan for pointing out a misunderstanding of the structure of L-parameters in [GS23], and professor A.-M. Aubert and Heiermann for correcting some imprecision in §5.1 related to the construction in [Hei06].
The author also sincerely thanks professor A.-M. Aubert for her careful reading of the draft and bringing related sections of [Hai14], [AX] to him.

### v3

@@ 7. An Application to $\mathrm{G}_{2}$
In this section, as an application of the main theorem, we verify the formal degree conjecture for discrete series of split $\mathrm{G}_{2}$ supported on a maximal Levi subgroup. Let $\mathbf{G}=\mathrm{G}_{2}$ be the split exceptional group of type $\mathrm{G}_{2}$ over $F$. We first recall the classification of discrete series representations of $G=\mathbf{G}(F)$ supported on a maximal Levi.
@@
For convenience of the reader, we list the related normalizations and the decomposition $r=\oplus_{i}r_{i}$ (the adjoint representation of ${{}^{L}\!M}$ on $\widehat{\mathfrak{n}}$) for these two cases as follows.
@@
    In this case the representation $r$ of ${}^{L}\!M_{\alpha}$ on $\widehat{\mathfrak{n}}_{\alpha}$ decomposes as
@@
    In this case the representation $r$ of ${}^{L}\!M_{\beta}$ on $\widehat{\mathfrak{n}}_{\beta}$ decomposes as
@@
The discrete series of $G$ has been classified by Muić. Here we just summarize below the cuspidal support of those supported on a maximal parabolic; see the original paper [muic-unitary-dual-of-g2] of Muić, and another convenient reference is [gan-savin-g2theta, §2].
@@
The full explicit local Langlands correspondence for $\mathrm{G}_{2}$ has been established in [gan-savin-g2-llc] based on a series work in exceptional theta correspondence. The conjectural L-parameters for non-supercuspidal discretes series have already been classified in [gan-savin-g2theta, §3.5], which we now recall.
@@
$\pi$ $\varphi(\mathrm{SL}_{2})$ $\varphi(\mathrm{W}_{F})$ $\pi_{0}(\overline{S}_{\varphi})=\mathcal{S}_{\varphi}^{\natural}$
@@
by [gan-savin-g2theta, §3.5]. (Note that the representations $\pi_{\alpha}(\tau,1/2)$, $\pi_{\beta}(\tau,1/2)$, $\pi_{\beta}(\tau,1)$ above are denoted by $\delta_{P}(\tau)$, $\delta_{Q}(\tau)$, $\pi_{\mathrm{gen}}[\tau]$, respectively, in [gan-savin-g2theta].)
@@
For the structure constants we have the following table:
$\mathbf{G}$ $\gamma$ $\mathbf{M}_{\gamma}$ $m$ $\chi\in X^{*}(\mathbf{M}_{\gamma})^{\mathbf{G}}$ generator $\gamma^{\vee}$ $\frac{m}{\langle\chi,\gamma^{\vee}\rangle}$
$\mathrm{G}_{2}$ long $\beta$ $\mathrm{GL}_{2}$ $2$ $\det=(3\alpha+2\beta)=\tilde{\beta}$ short $\beta^{\vee}$ $2$
$\mathrm{G}_{2}$ short $\alpha$ $\mathrm{GL}_{2}$ $2$ $\det=(2\alpha+\beta)=\tilde{\alpha}$ long $\alpha^{\vee}$ $2$
Since $\mathrm{G}_{2}$ is semisimple and adjoint, $\mathcal{S}_{\varphi_{\pi}}^{\natural}=\pi_{0}(\overline{S}_{\varphi_{\pi}})$, while for $\mathbf{M}_{\gamma}\simeq\mathrm{GL}_{2}$ we have $\mathcal{S}_{\varphi_{\tau}}^{\natural}=2$, so that the result of our main theorem reduces to
$\frac{|\mathcal{S}_{\varphi_{\pi}}^{\natural}|}{|\mathcal{S}_{\varphi_{\sigma}}^{\natural}|}=j\cdot\frac{\langle\chi,\alpha^{\vee}\rangle}{m}\Longleftrightarrow|\pi_{0}(\overline{S}_{\varphi_{\pi}})|=j,$
which is clearly compatible with the above three classes of L-parameters.
The compatibility of Langlands-Shahidi $\gamma$-factors and those from LLC for split $\mathrm{G}_{2}$ has been verified in [shahidi-g2]. Of course, the formal degree conjecture is known for $\mathbf{M}_{\alpha},\mathbf{M}_{\beta}\simeq\mathrm{GL}_{2}$; thus as an application we have verified the formal degree conjecture for discrete series of $\mathrm{G}_{2}$ supported on maximal Levi subgroups.
@@ Acknowledgements
The author would like to first thank his advisor A. Ichino for suggesting this problem and many helpful discussions. He thanks professor W. T. Gan for pointing out a misunderstanding of the structure of L-parameters in [gan-savin-g2theta], and professor A.-M. Aubert and V. Heiermann for correcting some imprecision in §5.1 related to the construction in [heiermann-orbites].
The author also sincerely thanks professor A.-M. Aubert for her careful reading of the draft and bringing related sections of [haines], [aubert-xu] to him.
Data Availability Statement
The author confirms that all data generated or analyzed in this study are available within the article.
[…]
[heading=bibintoc,]